OpenStudy (anonymous):

Use first principles to find the derivative of h(x) = 2sqrtx

OpenStudy (anonymous):

\[\lim _{ {h \rightarrow 0}} \frac{f(x+h)-f(x)}{h}\], if I'm not mistaken?

OpenStudy (anonymous):

@pingpong21 you there?

OpenStudy (anonymous):

Sorry, but yest that's correct

OpenStudy (anonymous):

:) Thanks! Need some help with that limit?

OpenStudy (anonymous):

Yeah, just how to get to the final answer

OpenStudy (anonymous):

To get the derivative I mean

OpenStudy (anonymous):

Sorry, I had some internet problems.... Could you get the answer in the meantime? or should I quickly help?

OpenStudy (anonymous):

Could use some help if you can

OpenStudy (anonymous):

No problem! Okay, so \[\lim _{ {h \rightarrow 0}} \frac{f(x+h)-f(x)}{h} \\ =\lim _{ {h \rightarrow 0}} \frac{2\sqrt{x+h}-2\sqrt{x}}{h}\] and that equals\[\lim _{ {h \rightarrow 0}} \frac{2}{\sqrt{x+h}+\sqrt{x}}\]

OpenStudy (anonymous):

Now, by the properties of limits, we get\[=\frac{2}{\color {red}{(\lim _{ {h \rightarrow 0}}\sqrt{x+h})}+\sqrt{x}}\]

OpenStudy (anonymous):

can you try to solve that?

OpenStudy (anonymous):

Wouldn't it just disappear? Given h goes to 0

OpenStudy (anonymous):

and what rule allows you to drop the sqrt to the denominator?

OpenStudy (anonymous):

Yes! h goes to 0, so you can replace it with 0.

OpenStudy (anonymous):

The fraction was rationalized, that is making it so that there are no sqrt's and stuff in the numerator

OpenStudy (anonymous):

alright, thank you very much

OpenStudy (anonymous):

No problem:) Looks like you understand what to do?

OpenStudy (anonymous):

Yup. Just drop the sqrt to the bottom if it shows

OpenStudy (anonymous):

An easy way is also to write the roots as exponentials (e.g sqrt(x) = x^(1/2)). However, in this specific case it wouldn't really have helped